23.) A 50.0-g Super Ball traveling at 30.0 m/s bounces off a brick wall and rebounds at 18.0 m/s. A high-speed camera records this event. If the ball is in contact with the wall for 4.10 ms, what is the magnitude of the average acceleration of the ball during this time interval? m/s2 24.) An artillery shell is fired with an initial velocity of 300 m/s at 54.5° above the horizontal. To clear an avalanche, it explodes on a mountainside 38.0 s after firing. What are the x- and y-coordinates of the shell where it explodes, relative to its firing point? x = m y = m
Answer #1

here,

mass , m = 50 g

initial speed , u = - 30 m/s

final speed , v = 18 m/s

time of contact , t = 0.0041 s

let the average accelration be a

v = u + a * t

18 = - 30 + a * 0.0041

a = 11707.32 m/s^2

a = 1.17 * 10^4 m/s^2

the average accelration is 1.17 * 10^4 m/s^2

b)

initial speed , u = 300 m/s

theta = 54.5 degree

t = 38 s

x = u * cos(theta) * t

x = 300 * cos(54.5) * 38

x = 6620.01 m

y = u * sin(theta) * t - 0.5 * g * t^2

y = 300 * sin(54.5) * 38 - 0.5 * 9.8 * 38^2

y = 2205.32 m